The results of a titration experiment were as follows:

 Volume of aqueous NaOH pipetted into the conical flask = 10.00 ml

 Average volume of dilute HCI added from the burette to reach the

endpoint = 12.54 mL.

 Molarity of the dilute HCI = 1.196 M


1- Calculate the number of moles of HCI added during the titration.

-What is the number of moles of NaOH that has reacted during the titration?

-What is the molarity of the aqueous NaOH?


1
Expert's answer
2021-04-29T07:26:03-0400

"HCl(aq) + NaOH(aq) \u2192 NaCl(aq) + H2O(l)"


Volume of HCl = 12.54-10.00=2.54ml


"Moles = 1.196×\frac{2.54}{1000}=3.04×10^{-3}Moles"



Moles of NaOH"= 3.04\u00d710^{-3}Moles"


Molarity of NaoH"\frac{3.04×10^{-3}}{10.00×10^{-3}}=0.304M"